OpenStudy (anonymous):

carry out the following multiplication (1.38 x 10^2)(2.58 x 10^-7)

OpenStudy (anonymous):

= 35604x10^-4

OpenStudy (unklerhaukus):

\[(1.38 \times 10^2)(2.58 \times 10^{-7})\\=(1.38\times2.58)\times10^{2-7}\\=3.5604\times10^{-5}\]